A monkey jumps from a branch that is 5 meters above the ground. Ignoring air resistance how long does it take to reach the ground?
1
Expert's answer
2020-12-08T10:46:34-0500

We can find the time that the monkey takes to reach the ground from the kinematics equation:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot 5\ m}{9.8\ \dfrac{m}{s^2}}}=1.01\ s."

Answer:

"t=1.01\ s."